OpenStudy (anonymous):

Let f(x) = x^2 + 1 lim h -> 0 [f(x+h) - f(x)]/h

OpenStudy (jhannybean):

\[\large \lim_{h \rightarrow 0} \frac{{(x+h)}-f(x)}{h}\] Given function \[\large f(x)=x^2+1\] plug your f(x) into your "x" value into the definition of a derivative.\[\large \lim_{h \rightarrow 0}\frac{[(x^2+1)+h]-(x^2+1)}{h}\] are you able to evaluate this now?

OpenStudy (luigi0210):

Made a little booboo Jhanny :P

OpenStudy (jhannybean):

Where at?

OpenStudy (jhannybean):

Oh i forgot the f in the definition?

OpenStudy (luigi0210):

(x+h)^2+1

OpenStudy (jhannybean):

ahh...yeah.

OpenStudy (anonymous):

2x

OpenStudy (jhannybean):

\[\large \lim_{h \rightarrow 0} \frac{f{(x+h)}-f(x)}{h}\] Given function \[\large f(x)=x^2+1\] plug your f(x) into your "x" value into the definition of a derivative.\[\large \lim_{h \rightarrow 0}\frac{[(x+h)^2+1]-[x^2-1]}{h}\] are you able to evaluate this now?

OpenStudy (jhannybean):

omg it'ssupposed to be x^2+1 inside. -_- i hate latex :|

OpenStudy (anonymous):

latex LOL

OpenStudy (jhannybean):

\[\large \lim_{h \rightarrow 0}\frac{[(x+h)^2+1]-[x^2+1]}{h}\]\[\large \lim_{h \rightarrow 0}\frac{x^2+2xh+h^2+1-x^2-1}{h}\]\[\large \lim_{h \rightarrow 0}\frac{2xh+h^2}{h}\]factor out an h. \[\large \lim_{h \rightarrow 0}\frac{\cancel h(2x+h)}{\cancel h}\] when h goes to 0, you get 2x as your answer.

OpenStudy (anonymous):

thanks

OpenStudy (jhannybean):

np, understand how it works?

OpenStudy (luigi0210):

nope :)

OpenStudy (anonymous):

yes